Can someone please help me solve PART D
Answer #1

D) When the battery is connected energy stored in the capacitor,

U2 = (1/2)*C*V^2

= (1/2)*(k*A*epsilon/d)*V^2

= (1/2)*(3*30*10^-4*8.854*10^-12/(5*10^-3))*12.5^2

= 1.24*10^-9 J

when the battery is disconnetced,

U3 = 1.66*10^-9 J


Workdone, W = decrease in potential energy

= U3 - U2

= 1.66*10^-9 - 1.24*10^-9

= 4.2*10^-10 J